OpenStudy (anonymous):

simplify: 72/5 + 94/5

OpenStudy (anonymous):

\[7\frac{ 2 }{ 5 }+\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

\[\frac{72}{5}+\frac{94}{5}\] If the denominator of 2 fractions are same, you can simply add the numerator as you normally would \[\implies \frac{72}{5}+\frac{94}{5}=\frac{(72+94)}{5}\] After addition if there are any common factors, they may be cancelled out

OpenStudy (anonymous):

thats makes no sense

OpenStudy (anonymous):

there is dno 72 i wrote it wrong